OpenStudy (anonymous):

find x intercept 2x+4y=24 i got a weird number so i want to double check if it is correct

OpenStudy (aroub):

what did you get? :)

OpenStudy (anonymous):

.5/6

OpenStudy (anonymous):

for y i got 6 which makes sense but i dont get the x

OpenStudy (aroub):

wait.. all you have to do to find the x-intercept is to make y=0 , how? , look, 2x+4y=24 2x+4(0)=24 2x=24 x=12 and for the y-intercept x=0 2x+4y=24 2(0)+4y=24 4y=24 y=6 get it?

OpenStudy (anonymous):

oh ok i get it, someone had taught me to re write first so i ended up with y=-1/2x+6

OpenStudy (aroub):

its right by the way BUT only you do that if you want to find the y-intercept ... :)

OpenStudy (anonymous):

oooo ok

OpenStudy (aroub):

=)